********PLEASE PLEASE PLEASE ANSWER 1-3****** THANK YOUU!!!

1a.When 22.0 mL of a 8.66×10-4 M cobalt(II) iodide solution is combined with 25.0 mL of a 1.92×10-4 M ammonium sulfidesolution does a precipitate form?   (yes or no)

For these conditions the Reaction Quotient, Q, is equal to______

1b.When 22.0 mL of a 6.52×10-4M ammonium iodide solution is combined with 12.0 mL of a 1.05×10-4M lead acetatesolution does a precipitate form?  (yes or no)

For these conditions the Reaction Quotient, Q, is equal to_____.

2a.A 26.2 mL sample of a 0.407 M aqueous acetic acid solution is titrated with a 0.392 M aqueous sodium hydroxide solution. What is the pH at the start of the titration, before any sodium hydroxide has been added?

pH = _____

2b.What is the pH at the equivalence point in the titration of a 22.0 mL sample of a 0.388 M aqueous hypochlorous acid solution with a 0.447 M aqueous barium hydroxide solution?

pH = ______

3a.How many grams of solid sodium fluoride should be added to 0.500 L of a 0.107 M hydrofluoric acid solution to prepare a buffer with a pH of 3.677 ?

grams sodium fluoride =______ g.

3b.How many grams of solid ammonium chloride should be added to 1.50 L of a 0.281 M ammonia solution to prepare a buffer with a pH of 9.990 ?  

grams ammonium chloride =_______ g.
Answer #1

Ans 1a

Moles of cobalt(II) iodide = molarity x volume

= 8.66*10^-4 mol/L * 22 mL * 1L/1000mL

= 0.000019052 mol

Moles of ammonium sulfide = molarity x volume

= 1.92*10^-4 mol/L * 25 mL * 1L/1000mL

= 0.0000048 mol

The balanced reaction

CoI2(aq) + (NH4)2S(aq) = 2NH4I(aq) + CoS(s)

Possible precipitate in this reaction

CoS(s) = Co2+(aq) + S2-(aq)

Reaction quotient

Q = [Co2+] [S2-]

After mixing

Volume = 22+25 = 47 mL = 0.047 L

Molarity of cobalt cation = 0.000019052 mol / 0.047 L

= 0.0004054 M

Molarity of sulfide anion = 0.0000048 mol / 0.047 L

= 0.00010213 M

Q = 0.0004054 * 0.00010213

= 4.14*10^-8

Actual Ksp value of CoS = 5.9*10^-21

Q > actual Ksp

Therefore a precipitate will form.